Question 5 (1 point)

Stanley Roper has $2,400 that he is looking to invest. His brother approached him with an investment opportunity that could give Patrick $4,600 in 4 years. What interest rate would the investment have to yield in order for Stanley’s brother to deliver on his promise? (Answer needs to be stated as a decimal. For example: .1192) Round to four decimal places.

Your Answer:

Question 5 options:

Answer

Question 6 (1 point)

Chuck Brown will receive from his investment cash flows of $3,155, $3,490, and $3,800 at the end of years 1, 2 and 3 respectively. If he can earn 7.5 percent on any investment that he makes, what is the future value of his investment cash flows at the end of three years? (Round to two decimal places.)

Your Answer:

Question 6 options:

Answer

Question 7 (1 point)

Your brother has asked you to help him with choosing an investment. He has $8,000 to invest today for a period of two years. You identify a bank CD that pays an interest rate of 0.0300 annually with the interest being paid quarterly. What will be the value of the investment in two years? Round to two decimal places.

Your Answer:

Question 7 options:

Answer

Question 8 (1 point)

You are evaluating a growing perpetuity product from a large financial services firm. The product promises an initial payment of $22,000 at the end of this year and subsequent payments that will thereafter grow at a rate of 0.02 annually. If you use a discount rate of 0.08 for investment products, what is the present value of this growing perpetuity? Round to two decimal places.

Your Answer:

Question 8 options:

Answer

Submit Quiz
Answer #1

Question 5

Present value = $2400

Time period = n = 4 years

Future value = FV = $4600

interest rate that investment will have to yield = r

Future value is calculated using the formula:

FV = PV*(1+r)n

4600 = 2400*(1+r)4

(1+r)4 = 4600/2400

(1+r) = (4600/2400)1/4

1+r = 1.17662114144118

r = 1.17662114144118 - 1

r = 0.17662114144118

Answer -> 0.1766